OpenStudy (anonymous):

The velocity of sound in air is given by the equation where v is the velocity in meters per second and t is the temperature in degrees Celsius. Find the temperature when the velocity of sound in air is 369 meters per second. Round to the nearest degree. A. 507ºC B. 6,535ºC C. 7,081ºC D. 67ºC

OpenStudy (anonymous):

\[v = 20 \sqrt{273 + t}\]

OpenStudy (anonymous):

\[369=20\sqrt{273+t}\]\[\frac{369}{20}=\sqrt{273+t}\]\[(\frac{369}{20})^2-273=t\]

OpenStudy (anonymous):

so which answer would it be?